OpenStudy (anonymous):

Solve log4 (y – 9) + log4 3 = log4 81

OpenStudy (anonymous):

have you tried anything at all ? :)

OpenStudy (anonymous):

... ok, assuming we have: \[\LARGE \log_4(y-9)+\log_4(3)=\log_4(81 )\] so.. \[\LARGE \log_4[(y-9)\cdot 3]=\log_4(81 )\] from here we take logs off ... and \[\LARGE (y-9)\cdot 3=81 \] \[\LARGE 3y-9\cdot 3=81 \] can you do it now ? :)

OpenStudy (anonymous):

@Stephanizer